OpenStudy (anonymous):

What is the geometric mean of 275 and 11?

OpenStudy (jamesj):

The geometric mean of two positive numbers, x, y, is \[ \sqrt{xy} \] Hence for your problem, the geometric mean of 275 and 11 is \[ \sqrt{275 \times 11} = ... \]

OpenStudy (anonymous):

Thank you very much!

OpenStudy (jamesj):

sure thing